OpenStudy (anonymous):

Consider the functions f(x) = 2(x − 3) and g(x) = 3x. What is the value of (f − g)(3)?

OpenStudy (ajprincess):

First find (f-g). (f-g)(x)=f(x)-g(x) =2(x-3)-3x Then substitute 3 for x to find (f-g)(3) Cn u do?

OpenStudy (ajprincess):

@orphtroll cn u do?

OpenStudy (waleed_imtiaz):

2x-6-3x=-x-6.. put x=3 u will get -9